OpenStudy (anonymous):

what is the integral of 9(1+cos)^2

OpenStudy (anonymous):

=int(9(1+2cos+cos^2)) =9int(1+2cosx+(cos2x+1)/2) =9(x+2sinx+(0.5sin2x+x)/2) =9(3/2x+2sinx+1/4sin2x)